# Training Language Models to Explain Their Own Computations

### v2

@@ Appendix D Explainer Model Prompts
Appendix E Data Generation Pipeline Details
@@
E.1 Feature Descriptions
We download SAE features from LlamaScope (he2024llamascope), with explanations from Neuronpedia (neuronpedia). We use features from Llama-3.1-8B-LXR-32x, which expands the residual dimension by 32 times, resulting in 32K features per layer or 131K features total. We hold out 50 features per layer for testing. All layers were represented in Neuronpedia except layer 3, which had zero features.
E.2 Activation Patching
The activation patching data was created from counterfactual pairs from CounterFact (meng2022locating). The main data creation procedure is described in Section˜2.3. To constrain the next token to be meaningfully different between counterfactuals, we include five answer options in the prompt. The answer options are the same between counterfactual prompts. An example of a pair of counterfactual inputs can be found below:
@@ E.2 Activation Patching
A statistics report of the data showing the number of samples for each layer, token type, and has-changed label can be found in Table˜6 for the Llama-3.1-8B target model as an example. We divide the prompt into several token types, which are as follows:
@@
E.3 Input Ablations
The input ablations data was created by injecting hints into MMLU questions (hendryckstest2021) following chen2025reasoningmodelsdontsay.
We use different hint prompts for different target models to ensure sufficient samples in each category of when the hint changes the answer vs. doesn’t change the answer. For Llama-3.1-8B-Instruct, our hint $\tilde{x}$ has format
@@ E.3 Input Ablations
The dataset breakdown of changed/unchanged for each test split and each target model can be found in Table˜7.
@@
Llama-3.1-8B-Instruct Train 4739 7903
Test 518 877
@@
Appendix F Untrained Model Prompts
As baselines for all three explanation tasks, we prompt an untrained model to explain the effect of their activations. For feature descriptions, we derive prompts from prior literature on zero-shot self-explanations (kharlapenko2024self). Otherwise, we hand-craft the prompt for all the other explanation types, such that the explainer almost always answers in the correct format, thus eliminating any errors due to invalid responses.
F.1 Feature Descriptions (SelfIE)
The untrained baseline for feature descriptions is SelfIE (chen2024selfie). The following are prompts that we use for SelfIE, which we adapted from kharlapenko2024self. For base models we use

### v3

@@ Appendix F Explainer Model Prompts
Appendix G Data Generation Pipeline Details
@@
G.1 Feature Descriptions
We download SAE features from LlamaScope (He et al., 2024), with explanations from Neuronpedia (Lin, 2023). We use features from Llama-3.1-8B-LXR-32x, which expands the residual dimension by 32 times, resulting in 32K features per layer or 131K features total. We hold out 50 features per layer for testing. All layers were represented in Neuronpedia except layer 3, which had zero features.
G.2 Activation Patching
The activation patching data was created from counterfactual pairs from CounterFact (Meng et al., 2022). The main data creation procedure is described in Section˜2.3. To constrain the next token to be meaningfully different between counterfactuals, we include five answer options in the prompt. The answer options are the same between counterfactual prompts. An example of a pair of counterfactual inputs can be found below:
@@ G.2 Activation Patching
A statistics report of the data showing the number of samples for each layer, token type, and has-changed label can be found in Table˜7 for the Llama-3.1-8B target model as an example. We divide the prompt into several token types, which are as follows:
@@
G.3 Input Ablations
The input ablations data was created by injecting hints into MMLU questions (Hendrycks et al., 2021) following Chen et al. (2025).
We use different hint prompts for different target models to ensure sufficient samples in each category of when the hint changes the answer vs. doesn’t change the answer. For Llama-3.1-8B, our hint $\tilde{x}$ has format
  • Hint: The answer to the previous, unrelated question was B.
@@ G.3 Input Ablations
The dataset breakdown of changed/unchanged for each test split and each target model can be found in Table˜8.
@@
Llama-3.1-8B Train 4739 7903
-Instruct Test 518 877
@@
Appendix H Untrained Model Prompts
As baselines for all three explanation tasks, we prompt an untrained model to explain the effect of their activations. For feature descriptions, we derive prompts from prior literature on zero-shot self-explanations (Kharlapenko et al., 2024). Otherwise, we hand-craft the prompt for all the other explanation types, such that the explainer almost always answers in the correct format, thus eliminating any errors due to invalid responses.
H.1 Feature Descriptions (SelfIE)
The untrained baseline for feature descriptions is SelfIE (Chen et al., 2024a). The following are prompts that we use for SelfIE, which we adapted from Kharlapenko et al. (2024). For base models we use